A corner lot that originally was square lost 185 meters squared of area when one of the adjacent streets was widened by 3 meters and the other was widened by 5 meters. Find the new dimensions of the lot.
Answer by ankor@dixie-net.com(22740)

A corner lot, that originally was square, lost 185 meters squared of area when one of the adjacent streets was widened by 3 meter and the other was widened by 5 meters.
Find the new dimensions of the lot.
:
Let x = original length of the square lot
the
x^2 = original area of the lot
:
(x-3) = new length
and
(x-5) = new width
:
Original area - new area = 185 sq/m
x^2 - (x-5)(x-3) = 185
FOIL
x^2 - (x^2 - 8x + 15) = 185
x^2 - x^2 + 8x - 15 = 185
8x = 185 + 15
x = 200/8
x = 25 meter is the original side of the lost
then
25 - 3 = 22 meters is then new length
25 - 5 = 20 meters is the new width
:
:
Check this by finding the areas
25*25 = 625 sq/m
22*20 = 440
-------------
dif: 185 sq/m